a river excrusion travels x miles upstream at 3 miles per hour and back to the same place at 5 miles per hour . The round trip takes 48 minutes. State , interms of x, how long the journey takes:
a) on the upstream leg;
b) on the downstream leg;
c) in total
d) form an equation in x and solve it to find how far upstream the excruison goes.

Answer by kietra(57)

(A) The total time the upstream journey would take is 3x.
(B) The total time the downstream leg would take is 5x.
(C) The total journey would take the upstream leg plus the downstream leg or 3x+5x.
(D)3x + 5x = 48 minutes
8x=48
x=6
So the upstream leg would take 3x = 3(6) = 18 minutes.
The downstream leg would take 5x = 5(6) = 30 minutes.
Now, to find the distance, use rate*time= distance.
So, the rate of the upstream leg was 3mph and the amount of time was 18 minutes.
3mph * 18 minutes We have to also convert units so 18 minutes and 3 miles per HOUR are in the same units. So, multiply by (1hr/60 minutes)
(3miles/hour) * (18 min) * (1hr/60 minutes)= 54/60 miles= 9/10 miles